I'm very confused!
Please help!
The Equations I know are:
Theta = arch length/radius
Angular velocity = theta/time
1. A wheel is rotating at an angular velocity of 1.2 pi radians/s while a point on the circumference of the wheel travels at 9.6 pi m in 10 seconds.
a) How many revolutions does the wheel make in 1 min?

Answer by edjones(7311)

60*1.2*pi*r=72*pi*r traveled in a minute.
2*pi*r=1 revolution
72*pi*r/2*pi*r=36 revolutions
.
Ed